Convert (3–√,−1) into polar coordinates (r,θ) so that r≥0 and 0≤θ<2π.

1
Expert's answer
2022-06-13T23:24:36-0400
"r^2=(-\sqrt{3})^2+(-1)^2=4""r\ge0=>r=\sqrt{4}=2"

Quadrant III


"\tan \theta=\dfrac{-1}{-\sqrt{3}}=\dfrac{1}{\sqrt{3}}""\theta=\pi+\tan^{-1}(\dfrac{1}{\sqrt{3}})=\dfrac{7\pi}{6}""z_1=2(\cos\dfrac{7\pi}{6}+i\sin\dfrac{7\pi}{6})"